# Adaptive LPD Radar Waveform Design with Generative Deep Learning

### v1

@@ VI-C Sensing Performance
We fine tune our generator for 3,000 iterations with both the cWGAN-GP loss and ambiguity loss (Eq. 15) to improve the ambiguity characteristics of the generated waveforms. We evaluate the effectiveness of the ambiguity loss term by computing the mainlobe width (MW), peak sidelobe level (PSL), and integrated Doppler ambiguity (IDA) of the generated waveforms conditioned on all RF background classes at $\{-10,0,10,30\}$ dB SNR. We compare these to the same metrics computed for the SIDLE waveforms. Because of the variable signal length of the SIDLE waveforms, we zero pad all waveforms to 11,000 time samples prior to computing their ambiguity function.
We find that the ambiguity loss improves the MW, PSL, and IDA of the generated waveforms. To understand this in the context of LPD, we evaluate each model’s detectability with a detection network trained on an equal number of SIDLE and generated examples (8800 per class). The resulting detectability is shown as a function of MW, PSL, and IDA in Fig. 10. The generators used in these figures were fine-tuned with a variable ambiguity loss strength $\eta$, target peak $\chi=1000$, and zero-Doppler weight $\gamma=1000$. Larger $\eta$ leads to improved MW, PSL, and IDA at the cost of greater detectability. However, generated waveforms with no ambiguity loss retain MW, PSL, and IDA values comparable or better than those of the SIDLE waveforms. A comparison of ambiguity examples from SIDLE and the LPD generator can be seen in Fig. 11.

### v2

@@ VI-C Sensing Performance
We fine tune our generator for 3,000 iterations with both the cWGAN-GP loss and ambiguity loss (Eq. 15) to improve the ambiguity characteristics of the generated waveforms. We evaluate the effectiveness of the ambiguity loss term by computing the mainlobe width (MW), peak sidelobe level (PSL), and integrated Doppler ambiguity (IDA) of the generated waveforms conditioned on all RF background classes at $\{-10,0,10,30\}$ dB SNR. We compare these to the same metrics computed for the baseline waveforms. Because of the variable signal length of the SIDLE waveforms, we zero pad all waveforms to 11,000 time samples prior to computing their ambiguity function.
We find that the ambiguity loss improves the MW, PSL, and IDA of the generated waveforms. To understand this in the context of LPD, we evaluate each model’s detectability with the ResNet-18 detector trained on an equal number of baseline and generated examples (8800 per class). The resulting detectability is shown as a function of MW, PSL, and IDA in Fig. 11. The generators used in these figures were fine-tuned with a variable ambiguity loss strength $\eta$, target peak $\chi=1000$, and zero-Doppler weight $\gamma=1000$. Larger $\eta$ leads to improved MW, PSL, and IDA at the cost of greater detectability. However, generated waveforms with no ambiguity loss retain MW, PSL, and IDA values comparable or better than those of the baseline waveforms, and match the values associated with the RF background. A comparison of ambiguity examples from baseline and the LPD generator can be seen in Fig. 12.
Figure 11: Tunable Sensing Performance. Shown are detectability vs sensing metrics for LPD generators that have been fine-tuned with variable ambiguity loss strengths (shades of green). They are compared to the baseline classes (orange), and a baseline generator without ambiguity loss fine-tuning (blue). Leftmost plot’s x-axis represents mainlobe width—lower implies better range resolution. Center plot’s x-axis represents peak sidelobe level—lower implies better target separation. Rightmost plot’s x-axis represents integrated Doppler ambiguity—lower implies less velocity ambiguity. Adjusting the ambiguity loss strength ($\eta$) allows one to control the trade-off between sensing ability and detectability. Increasing the ambiguity loss strength yields a less ambiguous waveform with narrower mainlobes, improved peak sidelobe level, and lower Doppler ambiguity at the cost of increased detectability. All values of ambiguity loss strength yield waveforms that offer comparable or improved sensing metrics compared to the baseline classes while being harder to detect.
@@ VII Discussion
Background Interference. Does mimicking the RF environment make our waveform more susceptible to interference at the receiver? To better understand the effects of interference on our waveforms, we re-run our ambiguity analysis with the RF background signal present in the return signal. We interfere our generated and baseline waveforms with the RF background at 0 dB SNR. We find our method degrades similarly to the baseline waveforms under interference. The interfered ambiguity metrics for our generated waveforms consist of an average MW of 23.1 samples (from 19.7), PSL of 0.38 (from 0.25), and IDA of 10984 (from 1576)—lower is better for all metrics. The interfered ambiguity metrics for the baseline waveforms follow a similar trend, with an average MW of 108.2 samples (from 111.1), PSL of 0.3 (from 0.19), and IDA of 42553 (from 2345). Overall, MW is not significantly affected, but PSL is increased and IDA is significantly increased. Notably, the vast majority of degradation in performance for our generated waveforms occurs when generated signals are conditioned on AM and FM modulations. As noted in the previous paragraph, waveforms conditioned on these modulations generally have inferior ambiguity characteristics when compared to other generated waveforms.
@@ VIII Conclusion
Our presented detection results for both FAM and ResNet-18 detectors utilized energy-normalized data, as this takes radar pulse width into account and ensures all waveforms have the same effective detection range. However, this leads to variable SNRs for all waveforms. Of equal interest from the perspective of detection are power-normalized results, which provides detection behavior at fixed SNRs.
We provide power-normalized detection results for both the ResNet-18 and FAM detectors in Fig. 13 and Fig. 14 respectively. Signals of interest are fixed to 0 dB SNR. In general, the FAM detector continues to perform poorly at low false alarm rates, though some classes are detected more easily. Notably, the ResNet-18 detector improves significantly at this fixed SNR, detecting all baseline classes at all false alarm rates at over $95\%$. Our generated waveforms continue to perform well in this power-normalized case, retaining a $22\%\pm 5.6\%$ detection rate.
In comparison to the energy-normalized results, these power-normalized results emphasize the importance of temporally spreading waveform energy. Our current framework fixes the pulse width of the generated waveform, but allowing the generator to increase that pulse width could yield further improved LPD characteristics.
Figure 13: Power-normalized ResNet-18 ROC Curve. Receiver operating characteristic (ROC) curves for the supervised binary detection network on power-normalized data. Baseline and generated waveforms are summed with the RF background at a fixed SNR of 0 dB. Our generated waveforms remain strictly harder to detect.
@@ Acknowledgment
Figure 14: Power-normalized FAM ROC Curve. Receiver operating characteristic (ROC) curves for the FFT Accumulation Method (FAM)-based cyclostationary detector on power-normalized data. Baseline and generated waveforms are summed with the RF background at a fixed SNR of 0 dB. Our generated waveforms closely mirror the random detection chance. This detector still fails for low false alarm rates.